# High-Q cavity coupled to a high permittivity dielectric resonator for sensing applications

### v2

@@ III Coupled STO and elliptical cavity
In practice, eigenmode information is obtained through S-parameter measurements. Figure 7 illustrates the S-parameters for an excitation scheme using two antennas (as shown in Fig. 4), with a puck offset of 50 mm. Two key parameters in the measurement are the frequency sensitivity of the modes to variations in $\varepsilon_{r}$ and the $Q$-factor of the modes. The quality factor can be calculated using a 3 dB method and is also related to phase variations in the $S_{1,2}$ curve; larger phase changes (i.e., a steeper phase derivative) correspond to a higher $Q$-factor. Using the phase derivative offers the benefit of not only analyzing the peaks but also assessing the sensitivity of the transmission zero. This notch, located between two resonance peaks in the S-parameters, corresponds to a point of minimal signal transmission.
@@
Figure 7(c) illustrates the phase derivatives of the PEC cavity at various values of $\varepsilon_{\mathrm{r}}$. The distribution or pattern of the phase derivative near resonance frequencies exhibits a Lorentzian shape. For modes near 1.3 GHz (corresponding to the elliptical cavity’s dominant mode), the phase derivative amplitude shows high values, indicating a large $Q$-factor. For the notch mode, the phase derivative peak remains nearly constant across different values of $\varepsilon_{r}$. For a normal conducting cavity, such as copper, the $S$-parameters are similar; however, the phase derivative amplitudes at the peaks are smaller, indicating a lower $Q$-factor.
@@
Figure 8 presents the frequencies of the two peaks in the $S_{1,2}$ curve, along with the notch between them. The resonance behavior as a function of $\varepsilon_{r}$ resembles eigenmode analysis, with level repulsion influenced by the $r_{\mathrm{offset}}$. The frequency of the notch shifts similarly to that of the single STO puck calculated from the semi-analytical formula Eq. 1. The frequency sensitivity of modes 1 and 2 is generally lower than that of the notch as shown in Fig. 8(b). Similar to the eigenmode analysis, for $\varepsilon_{r}$ below 230, mode 1 is the elliptical cavity-dominant mode, whereas above 230, it becomes the second mode. For the elliptical cavity-dominant mode, frequency sensitivity peaks at $\varepsilon_{r}=230$, although this is accompanied by a decrease in its phase derivative amplitude that is related to $Q$-factor value. However, the product of frequency sensitivity and the phase derivative for the two modes is equal to that of the notch for a PEC cavity (Fig. 8(c)), whereas for a copper cavity (Fig. 8(d)), this product for the notch exceeds that of the two modes. This coupling allows for a trade-off between higher frequency sensitivity and a higher $Q$-factor, depending on the $\varepsilon_{r}$ value.
@@
To assess the coupling between the two modes in practice, a test case was established by placing the STO puck inside an elliptical cavity. Excitation is provided through two long rods positioned at the end flanges, and the STO puck is secured at the center of the cavity with a radial offset of 19 mm, mounted on a ceramic rod (see Fig. 9). The resonance of the STO puck, observed at 1.16 GHz in Fig. 9(b), is consistent with the predictions of the simulations, corroborating the analysis reported in this section. In addition to the fundamental mode of the STO puck and the elliptical cavity, many other higher-order modes can also be excited in the resonators, which are partially reproduced by the simulated model (Fig. 9(c)). Depending on the application, one or more of these modes with different frequencies, field patterns, and frequency sensitivities could also be selected for mode coupling.
@@ IV Discussion
Unlike the power detection scheme, the fundamental limit in frequency measurements comes from the best achievable frequency stability of the measurement parts and thus it is not directly related to the Nyquist noise and the ambient temperature, although cryogenic cooling may significantly improve the detection limit by increasing the quality factors. By using superconducting cavities and sapphire, it is possible to achieve $Q$-factors in the $10^{9}$ range [44] and bandwidths of the order of 1 Hz at microwave frequencies [45, 46]. This, also considering that the resolution of the resonator linewidth is typically limited, by electronic noise, to about 1 ppm [44], explains the unprecedented sensitivity of the frequency measurement approach.
By conjugating the frequency sensitivity of the high-$Q$ cavity and the tunability introduced by coupling it with the STO resonator (Sec. III), we can envisage the possible use of the hybrid system in sensing applications. Although the specific characteristics of a particular sensor depends by its geometry and working conditions, we can anticipate the potentialities of the proposed hybrid system in this context. In particular, coupled dielectric resonators have already been proposed for the realization of sensitive particle and radiation bolometers at frequency around 10 GHz. In the bolometer detector described in Ref. [12, 47, 20, 21, 22], two resonators, realized using Sapphire and STO pucks, are positioned one on the top of the other and coupled by fringing fields. In the weak coupling regime, the maximum sensitivity to frequency occurs when the frequency of the two resonators exactly coincides and [47]
$\left.\frac{\mathrm{d}f(T)}{\mathrm{d}T}\right|_{\mathrm{max}}=\frac{A}{(W_{\mathrm{STO}})^{2}}\frac{\mathrm{d}f_{\mathrm{STO}}}{\mathrm{d}T},$ (2)
@@
being $A$ the coefficient that accounts for the coupling between the resonators modes and $W_{\mathrm{STO}}$ the unperturbed linewidth of STO [12]. The sensitivity of the bolometer is thus related to the temperature dependence of the frequency of the STO resonator ($\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$). From Eq. 1 we obtain
@@
being $\alpha$ related to the dimensions of the STO puck. Eq. 2 shows that the maximum sensitivity can be amplified by properly arranging the $A/(W_{\mathrm{STO}})^{2}$ factor.
We focus here on the possibility to exploit the coupled cavity-STO resonator system as a bolometer, that is by replacing the originally proposed sapphire resonator with an elliptical cavity. The main advantages of this scheme are the possibility to tune the coupling strength by varying the STO puck position within the cavity and, in particular, to achieve higher couplings than those due to fringing fields of dielectric resonators positioned at a short distance [12]. As shown by electromagnetic simulations reported in Sec. III, the frequency sensitivity is quantified as $\mathrm{d}f/\mathrm{d}\varepsilon_{\mathrm{r}}=2.8\leavevmode\nobreak\ \mathrm{MHz}$ for $\varepsilon_{\mathrm{r}}\approx 230$. How this number is translated in $\mathrm{d}f/\mathrm{d}T$ sensitivity depends by the dimensions of the resonator and by the specific working conditions, in particular by the temperature range.
The temperature-dependence of the STO resonator frequency is not trivial owing to the structural transition around 100 K, the ferroelectric phase transition at 51 K and the stabilization of the quantum paraelectric phase below 5 K [35, 37, 36]. Let’s first consider temperatures between 20 and 80 K, where the relative permittivity of STO can be approximated as $\varepsilon_{\mathrm{r}}\approx 10^{5}/T\approx 2000$. This range was discussed in the original proposal by Gallop et al. [12] and the frequency sensitivity quantified as $\mathrm{d}f(T)/\mathrm{d}T=Af_{0}/(2W^{2}_{STO}T)$. Considering the outcome of simulations reported in Sec. III, we expect that higher couplings enable the improvement of $\mathrm{d}f(T)/\mathrm{d}T$ if the sapphire resonator is replaced by the metal cavity. Although the $Q$-factor of the normal-conducting cavity is typically lower than that of sapphire resonator [48], simulations suggest that, to obtain the maximum sensitivity, the change of the notch frequency should be considered (Fig. 8(d)).
@@
We shall consider also the temperature range in which the cavity has superconducting properties. In particular, a Nb coated TESLA cavity is superconducting below 9.2 K where the $Q$-factor of the fundamental mode rises above $10^{9}$ [40]. Ideally, when the superconducting cavity is cooled below the critical temperature, the STO puck should be warm up to $T\approx 25\,\mathrm{K}$ to display the maximum value of $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$ [22]. This requirement, however, introduces further experimental complexity. Fig. 2 shows that between 1 and 0.4 K the frequency of the STO resonator alone decreases monotonically with temperature; from the fit we derive $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T=10\,\mathrm{kHz/K}$ at 0.8 K. This number is much smaller than what obtained at $T\approx 20-25\,\mathrm{K}$ [22], however it is expected to increase if small STO pucks are used (Eq. 3), as required, when the permittivity rises up to $\varepsilon_{r}\approx 30000$, to match the frequencies of the STO resonator and cavity. Additionally, the hybrid system can be tuned to increase the coupling strength and compensate, at least partially, lower $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$ values in this temperature range (Eq. 2). Given $\varepsilon_{\mathrm{r}}\approx 30000$, we expect that resonators having typical diameter in the range of 1 mm are needed to achieve a significant resonant coupling with the 1.3 GHz cavity. Considering that the heat capacity of the puck, which is directly related to the mass, plays a significant role in the system’s effectiveness as a bolometer [47], we expect that reducing the puck’s size increases sensitivity by amplifying the temperature rise per unit of absorbed energy, thereby enhancing the system’s detection capabilities.
@@
Given the very peculiar characteristic of STO and the dependence of the STO resonator frequency by external factors such as applied electric field or mechanical stress, we just mention that possible applications of the proposed system of coupled cavity and STO puck are not limited to the realization of bolometers. Simulations indicate that the two cavities can be tuned to reach the strong coupling regime (Fig. 5). As expected, at resonance the quality factor of the superconducting cavity progressively decreases as the coupling increases. However, we note that eigenmodes having $Q$-factor in the $10^{5}-10^{6}$ range can still be obtained by adjusting $r_{\mathrm{offset}}$ and the detuning of the natural frequencies of superconducting cavity and STO resonator. These values are corroborated by transmission spectra measured at low temperature (Fig. 2), which show that the $Q$-factor of the STO resonator is in the $10^{4}$ range. This value can potentially be further improved by using opportunely prepared STO pucks. Electromagnetic simulations also show that the change of the frequency of the superconducting cavity modes location, and of the corresponding phase, are amplified as an effect of the coupling with the STO resonator (Fig. 7). In this regard, the phase derivative of $S_{1,2}$ shows a particularly steep response.
@@
We also mention that preliminary tests were carried out to test the stability of the STO mode frequency. In the case of measurements running for hours, the observed drift of the resonance frequency, due to external factors and without any feedback control, was in the range of 50 Hz both for the STO puck measured at 0.1 K and for superconducting cavities measured at 1.7 K. These values are strongly reduced, of about one order of magnitude, for timescales of few seconds or minutes. By improving the experimental set up, and introducing a reference phase-locked loop (PLL) frequency stabilization, it is expected a significant reduction of the frequency drift, thus allowing to exploit phase measurements and further improve the system sensitivity. Additionally, the frequency tuning of the detection cavity can be ensured by the ferroelectricity of the STO and can be obtained by applying a DC electric field to the dielectric resonator. The expected tuning range is in the order of few percent.
@@ V Conclusions
To conclude, we have studied the hybrid system composed by a high-$Q$ TESLA-shaped elliptical cavity and STO resonator, and investigated the effect of parameters, such as STO permittivity, puck dimensions and position within the cavity, that govern the coupling between the electromagnetic modes. Finite-element simulations show that the hybrid system offers great versatility to tune the coupling strength and achieve the strong coupling regime. These results are supported by test measurements carried out at room temperature using a copper cavity and a STO puck, and by the low temperature characterization of the STO resonator, which shows resonant frequency of 0.1 GHz and $Q$-factor of 10000 at 0.16 K. The hybrid system show potential for the realization of microwave sensors in which the sensitivity of the STO puck to selected physical quantities is exploited as the active element, while frequency or phase measurements on the high-$Q$ cavity are used to efficiently detect such changes. In particular, the application of the hybrid system in bolometers exploiting the temperature dependence of the STO permittivity has been discussed. Our results are useful to design dedicated experiments at low temperature allowing the direct test of sensitivity and tunability of the proposed hybrid system in high precision frequency-domain measurements.
VI ACKNOWLEDGMENTS

### v3

@@ III Coupled STO and elliptical cavity
In practice, eigenmode information is obtained by measuring the S-parameters. Figure 7 illustrates the S-parameters for an excitation scheme using two antennas (as shown in Fig. 4), with a puck offset of 50 mm. Two key parameters in the measurement are the frequency sensitivity of the modes to variations in $\varepsilon_{r}$ and the $Q$ factor of the modes. The quality factor can be calculated using a 3 dB method and is also related to phase variations in the $S_{1,2}$ curve; larger phase changes (i.e., a steeper phase derivative) correspond to a higher $Q$ factor. Using the phase derivative offers the benefit of not only analyzing the peaks but also assessing the sensitivity of the transmission zero. This notch, located between two resonance peaks in the S-parameters, corresponds to a point of minimal signal transmission.
@@
Figure 7(c) illustrates the phase derivatives of the PEC cavity at various values of $\varepsilon_{\mathrm{r}}$. The distribution or pattern of the phase derivative near the resonance frequencies exhibits a Lorentzian shape. For modes near 1.3 GHz (corresponding to the elliptical cavity’s dominant mode), the phase derivative amplitude shows high values, indicating a large $Q$ factor. For the notch mode, the phase derivative peak remains nearly constant across different values of $\varepsilon_{r}$. For a normal conducting cavity, such as copper, the $S$-parameters are similar; however, the phase derivative amplitudes at the peaks are smaller, indicating a lower $Q$ factor.
@@
Figure 8 presents the frequencies of the two peaks in the $S_{1,2}$ curve, along with the notch between them. The resonance behavior as a function of $\varepsilon_{r}$ resembles eigenmode analysis, with level repulsion influenced by $r_{\mathrm{offset}}$. The frequency of the notch shifts similarly to that of the single STO puck calculated from the semi-analytical Eq. 1. The frequency sensitivity of modes 1 and 2 is generally lower than that of the notch, as shown in Fig. 8(b). Similarly to the eigenmode analysis, for $\varepsilon_{r}$ below 230, mode 1 is the elliptical cavity-dominant mode, whereas above 230, it becomes the second mode. For the elliptical cavity-dominant mode, frequency sensitivity peaks at $\varepsilon_{r}=230$, although this is accompanied by a decrease in its phase derivative amplitude that is related to $Q$ factor value. However, the product of frequency sensitivity and the phase derivative for the two modes is equal to that of the notch for a PEC cavity (Fig. 8(c)), whereas for a copper cavity (Fig. 8(d)), this product for the notch exceeds that of the two modes. This coupling allows for a trade-off between higher frequency sensitivity and a higher $Q$ factor, depending on the $\varepsilon_{r}$ value.
@@
To assess the coupling between the two modes in practice, a test case was established by placing the STO puck inside an elliptical cavity. Excitation is provided through two long rods placed at the end flanges, and the STO puck is secured at the center of the cavity with a radial offset of 19 mm, mounted on a ceramic rod (see Fig. 9). The resonance of the STO puck, observed at 1.16 GHz in Fig. 9(b), is consistent with the predictions of the simulations, corroborating the analysis reported in this section. In addition to the fundamental mode of the STO puck and the elliptical cavity, many other higher-order modes can also be excited in the resonators, which are partially reproduced by the simulated model (Fig. 9(c)). Depending on the application, one or more of these modes with different frequencies, field patterns, and frequency sensitivities could also be selected for mode coupling.
@@ IV Discussion
Unlike the power detection scheme, the fundamental limit in frequency measurements comes from the best achievable frequency stability of the measurement parts, and thus it is not directly related to Nyquist noise and ambient temperature, although cryogenic cooling may significantly improve the detection limit by increasing the quality factors. Using superconducting cavities and sapphire, it is possible to achieve $Q$ factors in the $10^{9}$ range [52] and bandwidths of the order of 1 Hz at microwave frequencies [53, 47]. This, also considering that the resolution of the resonator linewidth is typically limited, by electronic noise, to about 1 ppm [52], explains the unprecedented sensitivity of the frequency measurement approach.
In the coupled system, the high-$Q$ cavity can be used as a stable frequency reference enabling high-precision measurements, while the coupling with the STO resonator adds tunability and sensitivity to external physical quantities (Sect. III). Although the specific characteristics of a particular sensor depend on its geometry, as well as by temperature and working conditions, we can envisage the potential of the coupled system in sensing applications. The maximum sensitivity of the frequency of the coupled system with respect to temperature occurs at resonance and reads
$\left.\frac{\mathrm{d}f(T)}{\mathrm{d}T}\right|_{\mathrm{max}}=\beta\frac{\mathrm{d}f_{\mathrm{STO}}}{\mathrm{d}T},$ (2)
@@
being $\beta$ a coefficient related to the coupling strength and the quality factor of the modes [54, 13]. The sensitivity is thus related to the temperature dependence of the STO frequency ($\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$), and weighted by the coefficient $\beta$.
In the bolometer detector described in [13, 54, 24, 25, 26], two dielectric resonators, made using Sapphire and STO pucks, are coupled by fringing fields. In our experiment, the Sapphire resonator is replaced with the elliptical cavity that allows a better control of the coupling strength by varying the position of the STO puck within the cavity (Fig. 5). A simple expression of the maximum sensitivity of the coupled system (Eq. 2) can be obtained using the coupled-mode theory [18, 19] under resonant conditions. Considering the frequencies of mode 1 ($f_{1}$) and mode 2 ($f_{2}$) and their derivative with respect to temperature, $\mathrm{d}f_{\mathrm{1(2)}}/\mathrm{d}T=\beta_{\mathrm{1(2)}}(\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T)$, the $\beta_{\mathrm{1}}$ and $\beta_{\mathrm{2}}$ coefficients read (Appendix A)
$\beta_{\mathrm{1}}=\frac{\sqrt{1-|\kappa|}}{1+\frac{Q_{\mathrm{STO}}}{Q_{\mathrm{cav}}}}$ (3)
and
$\beta_{\mathrm{2}}=\frac{\sqrt{1+|\kappa|}}{1+\frac{Q_{\mathrm{STO}}}{Q_{\mathrm{cav}}}},$ (4)
where the quality factors $Q_{\mathrm{STO}}$ and $Q_{\mathrm{cav}}$ refer to, respectively, the unperturbed dielectric resonator and cavity, and the coupling coefficient $\kappa$ is the normalized overlap integral of the electric fields of the dielectric resonator and cavity [18]. We note that the temperature sensitivity is improved using mode 2 at high $\kappa$, provided that $Q_{\mathrm{cav}}\gg Q_{\mathrm{STO}}$ (Eq. 3, 4). From the simulated curves in Fig. 5(a) we derive $\kappa$ considering that, on resonance ($\epsilon_{r}=230$), $f_{1}=f_{\mathrm{cav}}\sqrt{1-\kappa}$ and $f_{2}=f_{\mathrm{cav}}\sqrt{1+\kappa}$ [19]. At $r_{\mathrm{offset}}$=10 mm and $r_{\mathrm{offset}}$=50 mm, the coupling coefficients are respectively $\kappa=0.006$ and $\kappa=0.03$, which give rise to $\beta_{2}\approx 1$. Higher values of $\beta_{2}$ can be obtained by further improving the coupling strength. Being $\kappa$ a normalized quantity, Eq. 4 suggests that the upper limit of $\beta_{2}$ is $\sqrt{2}$.
The frequency of the STO resonator is dependent on the permittivity $\varepsilon_{\mathrm{r}}$, which has a remarkable temperature dependence. From Eq. 1 we obtain
@@
being $\alpha$ related to the dimensions of the STO puck. As shown by electromagnetic simulations (Sect. III), $\mathrm{d}f/\mathrm{d}\varepsilon_{\mathrm{r}}=2.8\leavevmode\nobreak\ \mathrm{MHz}$ for $\varepsilon_{\mathrm{r}}\approx 230$. At lower temperature, STO displays the structural transition around 100 K, the ferroelectric phase transition at 51 K and the stabilization of the quantum paraelectric phase below 5 K [40, 42, 41]. Between 20 and 80 K the relative permittivity of STO can be approximated as $\varepsilon_{\mathrm{r}}\approx 10^{5}/T\approx 2000$ [13]. Although in this range the $Q$ factor of the normal conductor cavity is typically lower than that of the sapphire resonator [55], simulations suggest that, to obtain the maximum sensitivity, the change in the notch frequency should be considered (Fig. 8(d)). Below 9.2 K a Nb-coated TESLA cavity is superconducting and the $Q$ factor of the fundamental mode increases above $10^{9}$ [45]. In optimal conditions, the superconducting cavity is cooled below the critical temperature while the STO puck is heated to $T\approx 25\,\mathrm{K}$ to display the maximum $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T$ [26]. However, at temperatures between 1 and 0.4 K the frequency of the STO resonator decreases monotonically with temperature (Fig. 2). From the fit, we derive $\mathrm{d}f_{\mathrm{STO}}/\mathrm{d}T=-10\,\mathrm{kHz/K}$ at 0.8 K. This number is much smaller than what was obtained at $T\approx 20-25\,\mathrm{K}$ [26], yet it is expected to increase if small STO pucks are used (Eq. 5), as required, when the permittivity increases to $\varepsilon_{r}\approx 30000$, to match the frequencies of the STO resonator and cavity. Given $\varepsilon_{\mathrm{r}}\approx 30000$, we expect that resonators having a typical diameter in the range of 1 mm are needed to achieve a significant resonant coupling with the 1.3 GHz cavity. Considering also that the heat capacity of the puck, which is directly related to the mass, plays a significant role in the effectiveness of the system as a bolometer [54], we expect that reducing the size of the puck increases sensitivity by amplifying the temperature increase per unit of absorbed energy, thus enhancing the detection capabilities of the system.
@@
Given the very peculiar characteristic of STO and the dependence of the STO resonator frequency by external factors such as applied electric field or mechanical stress, we just mention that possible applications of the proposed system of coupled cavity and STO puck are not limited to the realization of bolometers. Simulations indicate that the two cavities can be tuned to reach the strong coupling regime (Fig. 5). As expected, at resonance, the quality factor of the superconducting cavity progressively decreases as the coupling increases. However, we note that eigenmodes that have a $Q$ factor in the $10^{5}-10^{6}$ range can still be obtained by adjusting $r_{\mathrm{offset}}$ and the detuning of the natural frequencies of the superconducting cavity and the STO resonator. These values are corroborated by transmission spectra measured at low temperature (Fig. 2), which show that the $Q$ factor of the STO resonator is in the $10^{4}$ range. This value can potentially be further improved by using opportunely prepared STO pucks. Electromagnetic simulations also show that the change in the frequency of the superconducting cavity mode and of the corresponding phase is amplified as an effect of the coupling with the STO resonator (Fig. 7). In this regard, the phase derivative of $S_{1,2}$ shows a particularly steep response.
@@
We also mention that preliminary tests were carried out to test the stability of the STO mode frequency. In the case of measurements running for hours, the observed drift of the resonance frequency, due to external factors and without any feedback control, was in the range of 50 Hz both for the STO puck measured at 0.1 K and for superconducting cavities measured at 1.7 K. These values are strongly reduced, of about one order of magnitude, for timescales of a few seconds or minutes. By improving the experimental setup, and introducing a reference phase-locked loop (PLL) frequency stabilization, a significant reduction of the frequency drift is expected, thus allowing one to exploit phase measurements and further improve the system sensitivity. Additionally, the frequency tuning of the detection cavity can be ensured by the ferroelectricity of the STO and can be obtained by applying a DC electric field to the dielectric resonator. The expected tuning range is on the order of a few percent.
@@ V Conclusions
To conclude, we have studied the hybrid system composed by a high-$Q$ TESLA-shaped elliptical cavity and STO resonator, and investigated the effect of parameters, such as STO permittivity, puck dimensions and position within the cavity, that govern the coupling between the electromagnetic modes. Finite element simulations show that the hybrid system offers great versatility in tuning the coupling strength and achieving the strong coupling regime. These results are supported by test measurements carried out at room temperature using a copper cavity and a STO puck, and by the low-temperature characterization of the STO resonator, which shows a resonant frequency of 0.1 GHz and $Q$ factor of 10000 at 160 mK. The hybrid system shows potential for the realization of microwave sensors in which the sensitivity of the STO puck to selected physical quantities is exploited as the active element, while frequency or phase measurements in the high-$Q$ cavity are used to efficiently detect such changes. Our results are useful in designing dedicated experiments at low temperature, allowing the direct test of sensitivity and tunability of the proposed hybrid system in high-precision frequency-domain measurements.
Acknowledgements.
@@
Appendix A Temperature sensitivity of the coupled system
We apply the coupled-mode theory (CMT) [18, 19] to derive a simple expression for the sensitivity of the coupled cavity-STO resonator system as a function of temperature. The components of the electromagnetic field are expressed as a linear superposition of the uncoupled ones, i.e. the electric field of the coupled system is $\mathbf{E}=a_{1}\mathbf{E}_{\mathrm{STO}}+a_{2}\mathbf{E}_{\mathrm{cav}}$, being $\mathbf{E}_{\mathrm{STO}}$ the electric field of the dielectric resonator, $\mathbf{E}_{\mathrm{cav}}$ the electric field of the cavity and $a_{1},a_{2}$ the coupling coefficients. Similarly, the magnetic field is $\mathbf{H}=b_{1}\mathbf{H}_{\mathrm{STO}}+b_{2}\mathbf{H}_{\mathrm{cav}}$.
We assume that the fields are normalized ($\int_{V}\epsilon\mathbf{E}_{\mathrm{STO}}\cdot\mathbf{E}_{\mathrm{STO}}dV=\int_{V}\epsilon\mathbf{E}_{\mathrm{cav}}\cdot\mathbf{E}_{\mathrm{cav}}dV=A$), and consider the frequency ($f_{1}$) and quality factor ($Q_{1}$) of mode 1, and the frequency ($f_{2}$) and quality factor ($Q_{2}$) of mode 2. The ratio between coupled mode frequency and quality factor is [19]
$\frac{f_{1(2)}}{Q_{1(2)}}=\frac{|a_{1}|^{2}f_{\mathrm{STO}}}{Q_{\mathrm{STO}}}+\frac{|b_{\mathrm{2}}|^{2}f_{\mathrm{cav}}}{Q_{\mathrm{cav}}}$ (6)
where $f_{\mathrm{STO}}$ and $Q_{\mathrm{STO}}$ are respectively the frequency and quality factor of the uncoupled dielectric resonator, while $f_{\mathrm{cav}}$ and $Q_{\mathrm{cav}}$ are the frequency and quality factor of the uncoupled cavity, respectively.
We also assume that the frequency of the dielectric resonator depends on temperature ($f_{\mathrm{STO}}=f_{\mathrm{STO}}(T)$), while the quality factors are, to the first approximation, temperature independent at least in a small temperature interval around the working point. The derivative with respect to temperature is
$\frac{\mathrm{d}f_{1(2)}}{\mathrm{d}T}=|a_{1}|^{2}\frac{Q_{1(2)}}{Q_{\mathrm{STO}}}\frac{\mathrm{d}f_{\mathrm{STO}}(T)}{\mathrm{d}T};$ (7)
on resonance ($f_{\mathrm{STO}}=f_{\mathrm{cav}}$), the coefficients are $|a_{1}|=|a_{2}|=|b_{1}|=|b_{2}|=1/\sqrt{2}$ [19].
We consider the nonzero coupling coefficient $\kappa=\frac{\epsilon_{0}(\epsilon_{r}-1)}{A}\int_{V_{\mathrm{STO}}}\mathbf{E}_{\mathrm{STO}}^{*}\cdot\mathbf{E}_{\mathrm{cav}}dV$, where $V_{\mathrm{STO}}$ is the volume of the dielectric resonator [18]. On resonance, the frequency and quality factor of the coupled modes are [19]: $f_{1}=f_{\mathrm{cav}}\sqrt{1-\kappa}$ and
$Q_{1}=2\sqrt{1-|\kappa|}\frac{Q_{\mathrm{STO}}Q_{\mathrm{cav}}}{Q_{\mathrm{STO}}+Q_{\mathrm{cav}}}$ (8)
for mode 1; $f_{2}=f_{\mathrm{cav}}\sqrt{1+\kappa}$ and
$Q_{2}=2\sqrt{1+|\kappa|}\frac{Q_{\mathrm{STO}}Q_{\mathrm{cav}}}{Q_{\mathrm{STO}}+Q_{\mathrm{cav}}}$ (9)
for mode 2. The derivative with respect to temperature thus results
$\frac{\mathrm{d}f_{1}}{\mathrm{d}T}=\frac{\sqrt{1-|\kappa|}}{1+\frac{Q_{\mathrm{STO}}}{Q_{\mathrm{cav}}}}\frac{\mathrm{d}f_{\mathrm{STO}}}{\mathrm{d}T}$ (10)
for mode 1, and
$\frac{\mathrm{d}f_{2}}{\mathrm{d}T}=\frac{\sqrt{1+|\kappa|}}{1+\frac{Q_{\mathrm{STO}}}{Q_{\mathrm{cav}}}}\frac{\mathrm{d}f_{\mathrm{STO}}}{\mathrm{d}T}$ (11)
for mode 2.